4- 7d = 13 -4d


four minus 7d= thirteen minus 4 d

Found 2 solutions by prabhjyot, JimMarshall:
Answer by prabhjyot(165)

4- 7d = 13 -4d
4-13= -4d +7d
9 = 3d
d= 9/3=3
d=3

Answer by JimMarshall(82)

4- 7d = 13 -4d
==>4-13= -4d +7d
==>9 = 3d
==?d= 9/3=3
==>d=3
Ans